Fermi National Laboratory


Accelerator Update

Friday June 11
The day shift began with Operations monitoring store 3567, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:18 AM, the Main Injector (MI) tripped off due to a ground fault indication on the QD bus. Operations turned off MI and prepared the area for access.

Operations reported at 9:24 AM that Pbar experts began studies shortly after MI went down.

Operations reported at 11:05 AM that the Collider Detector at Fermilab (CDF) had not been able to take data from the present store due to colliding beam problems. Experts completed making collimator adjustments at 11:31 AM that allowed CDF to turn on.

MI experts reported at 12:11 PM that they found a bad quadrupole magnet (Q427). They began preparing to replace the magnet.

A Pbar expert began a study at 4:19 PM.

Booster experts finished a study period at 4:23 PM.

The Pbar expert completed his study at 5:51 PM. Operators prepared the Pbar tunnels for a two-hour access.

Operations reported at 11:14 PM that a lightning strike caused some Tevatron (TeV) orbits to move, but the store remained.

Saturday June 12
The midnight shift began with Operations monitoring store 3567, with MI off due to the quadruple magnet change, and with MiniBooNE taking beam.

At 1:11 AM, store 3567 aborted due to a RF problem (TRF4 and TRF8 tripped off) that occurred during a lightning strike.

Technicians in the MI tunnel reported at 2:27 AM that Q427 was successfully valved in to the MI LCW system.

Operations completed the MI highpot at 3:25 AM and it looked good.

Operations reported at 4:07 AM that a special beam detector for MiniBooNE got soaked during the storm and tripped off. Technicians will wait until daylight hours to make repairs. This held off MiniBooNE.

Operations established beam to MI at 4:36 AM.

TeV experts reported at 4:45 AM that they needed to work on a TeV RF station (TRF1).

Pbar resumed stacking at 4:53 AM.

Operations prepared the TeV for an access at 6:47 AM to make sector B3 vacuum pump repairs.

CDF began a two-hour access at 7:31 AM. D-zero (D0) began their access at 8:07 AM.

Techs completed their work on the MiniBooNE detector and MiniBooNE resumed taking beam at 8:11 AM.

A Recycler expert began a study period around shift change.

CDF came out of access at 10 AM and D0 came out at 10:30 AM.

Operations began a TeV checkout at 10:08 AM.

Operations reported at 10:38 AM that MiniBooNE was running at ten turns at 4.3 Hz for a total of 6.6E16 per hour.

High voltage experts brought feeder 23 back on line at 10:58 AM.

Operation started tuning beam to the Meson MCenter experiment at 11:08 AM.

At 11:52 AM, Operations reported that they had ramped the TeV to 980 GeV, but on the second ramp the TeV tripped off due to a sector E1 lead failure. The cryo system operators at CHL believed the problem was due to a readback error. A cryo system tech investigated.

At 2:28 PM, Booster tripped off due to a bad controller card for CEPSE1. Operators replaced the card and Pbar resumed stacking at 3:19 PM.

The cryo system tech completed his work on the E1 IO crate at 3:16 PM. Unfortunately a cold compressor tripped off. The cryo system pumpdown began at 3:32 PM. Operations put the TeV through a 150 GeV ramp at 4:46 PM.

At 5:51 PM, Operation tried to put some protons into the TeV, but found that the Tevatron Electron Lens (TEL) wasn't functioning properly. Operators searched for an expert to come in and fix the problem.

At 10:03 PM, a TEL expert discovered that a power supply, which had been replaced, had the wrong polarity.

Operations put the TeV into shot setup at 11:06 PM. Extraction of beam to Switchyard (SY) was disabled.

Sunday June 13
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 3571 at 1:22 AM with an initial luminosity of 68.04E30.

Operators had to make collimator adjustments and by 1:59 AM both CDF and D0 were taking data.

At 3 AM, Operations had to make more collimator adjustments to keep the colliding beam right.

Operation reestablished beam to MCenter at 3:30 AM.

At 8:51 AM, halted stacking and the MI ramp so that Recycler experts could conduct a one-hour study.

Pbar resumed stacking at 10:51 AM.

Meson experiment MTest requested beam at 10:45 AM.

At 5:46 PM, Operations reported that Booster had passed a Division Head's goal by extracting 6.03E12 protons in one pulse. The goal was 6.0E12.

At 9:34 a Booster kicker (MKS01) tripped off. An expert came in and had a bad cable repaired by 11 PM.

Monday June 14
The midnight shift began with Operations monitoring store 3571, with Pbar conducting studies, and with MiniBooNE, MTest, and MCenter taking beam.

Pbar experts completed their reverse proton study at 2:58 AM.

MI experts began a study period at 4:52 AM.

Plans for Today
The plan for today is to terminate the present store around 9 AM and put the TeV into shot setup around 1 PM today. TeV experts will conduct studies during the interim.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/14/2004   email Fermilab